OpenStudy (anikate):

you buy 6 bags of wild bird food to fill the feeders in your yard. Oyster shell grit sells for $4 a bag and sunflower seeds sell for $4.45 a bag. If you spend $25.80, how many of each type of feed are you buying?

OpenStudy (pfenn1):

If x is the number of bags of oyster shell grit then (6-x) is the number of bags of sunflower seeds. \[4x+4.45(6-x)=25.80\]

OpenStudy (anikate):

huh? I dont get it?

OpenStudy (pfenn1):

Okay another way to set it up. x=number of bags of oyster shell grit bought y=number of bags of sunflowers seeds bought You bought 6 bags of food all together so you know that x+y=6 y=6-x You also know that you spent altogether on the 6 bags is $25.80 The amount you spent on the bags of oyster seed grit would be the number of bags bought, x, multiplied by the cost per bag, which is $4.00=4.00x Likewise for the bags of sunflowers seeds $4.45y=4.45(6-x) So the total cost will be 4x + 4.45(6-x) = 25.80